OpenStudy (anonymous):

Find the zeros of the polynomial function with equation f(x) = x3-x2 -30x x = 0, x = 4, x = 5 x = 4, x = 5 x = -6, x = 5 x = 0, x = 6, x = -5

OpenStudy (anonymous):

factor this first. f(x)=x(x^2-x-30) f(x)=x(x-6)(x+5) Now solve each individual factor for zero to get the answers: 1. x=0 2. x-6=0 3. x+5=0

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

lol all of it..

OpenStudy (anonymous):

What is a zero, in terms of the graph of the function?